Two particles, with identical positive charges and a separation of 3.68 × 10-2 m, are released from rest. Immediately after the release, particle 1 has an acceleration a1 whose magnitude is 5.37 × 103 m/s2, while particle 2 has an acceleration a2 whose magnitude is 8.89 × 103 m/s2. Particle 1 has a mass of 7.43 × 10-6 kg. Find (a) the charge on each particle and (b) the mass of particle 2.
Answer #1

force on particle a = ma = 7.43*10^-6*5370 = 0.0425 N

force due to the two charges = kq1q2 / r^2

9*10^9*q1q2 / (0.0368)^2 = 0.0425

q1q2 = 6.4*10^-15

since charge is identical

q^2 = 8*10^-8C

mass of particle 2 = F/a = 0.0425 / 8890

= 4.78*10^-6 kg